Operation Summary for June 10, 2011 - June 13, 2011 |
|
The Integrated Luminosity for the period from 6/6/11 to 6/13/11 was 59.71 pb-1 (inverse picobarns). NuMI reported receiving 1.2E18 protons on target during this same period. Friday June 10 The day shift began with Operations monitoring the Tevatron (TEV) store 8804 (169.86E30) and the Recycler stash (44.0E10), with the Antiproton Source (Pbar) stacking (12.8E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations terminated store 8804 at 11:20 AM. Cryo system technicians worked on TEV sector C3 cryo system. Operations put the TEV into HEP shot setup and established store 8805 at 1:30 PM with an average initial luminosity (AIL) of 328.31E30. Operators reported that MiniBooNE was off from 12:43 PM to 4 PM for ground water work at MI-12. During the evening shift, Operations disabled beam to T-978 at 6:06 PM. Operators reported that Lambertson (LAM60) problems delayed the low intensity NuMI target scans; the scans started at 6:40 PM. Store 8805 aborted at 7:52 PM due to a corrector magnet power supply failure (HC49). Operators began the TEV turn-on and checkout. Operations put the TEV into shot setup and established store 8808 at 11:45 PM with an AIL of 271.59E30. Saturday June 11 The midnight shift began with Operations monitoring store 8808 (230.49E30) and the stash (49.4E10), with Pbar stacking (12.8E10), with NuMI conducting target scans, and with MiniBooNE taking beam. During the midnight shift, Operations reported that the NuMI target scans finished at 12:31 AM. Operations established beam to T-978 at 4:15 AM. Operators prepared MI-65 for access for a final safety walk-through. During the day shift, Operations established beam to NuMI at 10 AM and tuned for high intensity running. Operations terminated store 8808 at 3:16 PM and put the TEV into shot setup. During the evening shift, Operations established store 8809 at 4:26 PM with an AIL of 372.95E30. Operations disabled beam to T-978 at 6:04 PM. Sunday June 12 The midnight shift began with Operations monitoring store 8809 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to T-978 at 4:26 AM. At 7:27 AM, store 8809 aborted due to a bad vacuum indication at B3; the store was about to be terminated. During the day shift, operators investigated the abort, but could not find a problem. Operations put the TEV into shot setup and established store 8810 at 9:04 AM with an AIL of 379.65E30. During the evening shift, at 5:10 PM, operators reported a beam instability in Booster that resulted in bad beam positions. Operations turned off beam to MiniBooNE and decreased the beam intensity to NuMI and Pbar. Operations disabled beam to T-978 at 6:06 PM. Booster personnel came in and helped operators with the instability issues; together they were able to resume sending nominal intensities to Pbar, NuMI, and MiniBooNE at 8:49 PM. Monday June 13 The midnight shift began with Operations monitoring store 8810, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8810 at 2:30 AM, put the TEV into shot setup, and established store 8811 at 3:44 AM with an AIL of 403.48E30. Operations established beam to T-978 at 5:13 AM. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 06/13/2011 email Fermilab |
| Security, Privacy, Legal | |